1. Having covered the path from one plate to the other in a plane capacitor, an electron acquires a velocity of 108 cm/s. The distance between the plates is 5.3 mm. Find the difference of potentials between the plates.


1
Expert's answer
2022-01-30T13:52:58-0500

By conservation of energy:


"\dfrac12mv^2=\dfrac{1}{2}CV^2,\\\space\\
V=v\sqrt{\dfrac mC}=v\sqrt{\dfrac {md}{\epsilon_0}}=2.34·10^{-5}\text{ V}."